OpenStudy (anonymous):

Which property states that the order of numbers can be changed in an addition or multiplication problem? A. Associative B. Commutative C. Transitive D. Reflexive

OpenStudy (anonymous):

B.commutative as, a+b = b+a and axb = bxa